OpenStudy (anonymous):

Need help with adding subtracting decimals. In this unit, you learned about fractions that can be written as terminating decimals and fractions that can be written as nonterminating decimals. Create an addition word problem and a subtraction word problem using fractions. Show how to find the solution to each problem. Once you find the solutions to your problems, rewrite the solutions as either a terminating or nonterminating decimal.

OpenStudy (anonymous):

So you write a simple fraction addition problem: 1/2 + 1/5 and a subtraction problem: 1/2 - 1/5 Then you solve them, using what you learned this unit: 1/2+1/5=5/10+2/10=7/10 and 1/2-1/5=5/10-2/10=3/10 And then you convert them into decimals: 7/10 = 7 divided by 10 = 0.7, and 3/10 = 0.3 To get a nonterminating decimal, you use fractions that will result in a fraction such as 22/7, who's answer does not terminate